A manufacturer of graphing calculators has determined that 9,000 calculators per week will be sold at a price of $85 per calculator. At a price of $80, it is estimated that 11,000 calculators will be sold.

1. Determine a linear function that predicts the number of calculators that will be sold per week at a price of x dollars.

y = ____

2. Use this model to predict the number of calculators that will be sold at a price of $65.
1
Expert's answer
2020-09-30T20:00:11-0400

We know that,


"Slope=" change in y/change in x


If we represent number of calculators sold on y and decrease cost of calculators on the x axis.


So, the slope of line passing through the points (85, 9000) and ( 80, 11000)

"(x_1, y_1)" "(x_2, y_2)"

is.


"Slope=" change in number of calculator sold"\div" decrease in the cost of calculator


"Slope (m) =(y_2-y_1)\/(x_2-x_1)"


"Slope(m) =(11,000-9,000)\/(80-85)"


"Slope(m)=-400"


Therefore

"y-y_1=m(x-x_1)"

"m=-400"

"y_1=9,000"

"x_1=85"


"(y-9,000)=-400(x-85)"

"y-9,000=-400x+34,000"


"y=-400x+43,000"

Hence, the linear function the perdicts the number of calculators per week is given by "y=-400x+43,000"


2.

The number of calculators that will be sold at a price of $65.

"y=-400\u00d765+43,000"

"y=-26,000+43,000"

"y=17,000"